# Spatial correlations in SIS processes on regular random graphs

### v1

@@ III Dependence of nearest-neighbor correlation on network topology
We now further motivate the development of bespoke corrections to mean-field theory for the RRG by using numerical simulations to highlight significant variations in $F_{II}^{(1)}$, and hence global infection dynamics, with the underlying topology of regular networks. In order to do this, we investigate the effects of rewiring the regular lattice in two and three dimensions by randomly exchanging edges on the network with probability $p$, inspired by the methodology of Watts and Strogatz [14]. Here, we construct networks that preserve the degree of nodes, but become increasingly random in structure, in the sense that the probability of two adjacent nodes being connected by an edge decreases as $p$ increases. The edge swapping algorithm we use is a modification of an algorithm from Wang et al. [15] and can be summarized as follows:
  • Consider a network $G(V,E)$ representing a regular lattice, where $E$ and $V$ are the edge and node sets.
  • If $p_{e}\leq p$ swap the edges $e_{2i-1}=(u_{2i-1},v_{2i-1})$ and $e_{2i}=(u_{2i},v_{2i})$ to $e_{2i-1}^{\prime}=(u_{2i-1},v_{2i})$ and $e_{2i}^{\prime}=(u_{2i},v_{2i-1})$; Otherwise, keep the edges as they are, $e_{2i-1}^{\prime}=e_{2i-1}$ and $e_{2i}^{\prime}=e_{2i}$.
  • The revised network is $G=(V,E^{\prime})$ with $E^{\prime}=\{e_{1}^{\prime},e_{2}^{\prime},\dots,e_{|E|}^{\prime}\}$.
This algorithm yields randomized network structures that converge to the RRG as $p$ becomes ${\cal O}(1)$.
@@
In studying how increasingly random connectivity on the network affects spatial correlations between adjacent nodes, we first conduct a parameter sweep over the total rate of infection for each node, $k_{0}\beta$, and the probability of edge swapping, $p$, on the two- and three-dimensional square lattices, for which the node degree is $k_{0}=4$ and $k_{0}=6$, respectively. Varying $k_{0}\beta$ and $p$, we find that both an increase in $p$ (and hence, a reduced mean shortest path length on the network) and a higher rate of infection per node, $k_{0}\beta$, suppress spatial correlations between infectious nodes. This is manifested in shorter relaxation times (approximated by the time between peak and half-maximum values of $F_{II}^{(1)}(t)$) for lower mean-shortest path lengths and higher rates of infection (Figs. 3(a) and 3(c)). Furthermore, the steady-state value of $F_{II}^{(1)}\left(t\right)$ (Figs. 3(b) and 3(d)), which is in turn related to the steady-state infection density on the network, is also highly sensitive to both $p$ and the rate of infection. Similar effects have previously been observed on the proportion of recovered individuals in the closely-related SIR model [16, 17].
@@
Figure 3: Relaxation time and steady state values for the SIS process ($\gamma=1,N=10^{4},I_{0}=10^{-2}$) versus $k_{0}\beta$ and $p$. Figures (a) and (c) show log relaxation times, which we define as the time between the peak ($t_{\rm{peak}}$) and half-maximum ($t_{1/2}$) values of $F_{II}^{(1)}(t)$, whilst (b) and (d) show the value of $\ln(\bar{F}_{II}^{(1)}-1)$ for $k_{0}=4$ and $6$, respectively. The results of the parameter sweep show that increasing $p$ and increasing $k_{0}\beta$ both reduce the steady state value of $F_{II}^{(1)}\left(t\right)$ and the relaxation time for the process.
From the results of the parameter sweep, we see that changes in both $p$ and $k_{0}\beta$ strongly affect $F_{II}^{(1)}\left(t\right)$, and hence, the trajectory of infection density on the network as a function of time (via Eq. (6)). In general, adding topological randomness to the network reduces the relaxation time of the process and increases the steady-state infection value. As such, it is analogous to increasing $\beta$ in terms of its effect on spatial correlations between infectious nodes. The results of these simulations suggest that the corrections to mean-field theory proposed and analyzed in previous work on the regular lattice, where $p=0$ [11, 12], are inappropriate corrections for equivalent simulations on the RRG, where $p={\cal O}(1)$, due to a strong relationship between $p$ and $F_{II}^{(1)}\left(t\right)$. As such, in the proceeding section, we derive a novel correction to mean-field theory based on the theory of random networks, finding excellent agreement with corresponding numerical simulations that outperform both mean-field theory and previous corrections for the regular lattice.
@@ IV Dynamics of the Pairwise Correlation Function
We now formulate generalized dynamics for the distance-one pairwise correlation function, $F_{II}^{(1)}\left(t\right)$, for both the lattice and the RRG, using a series of hierarchical ODEs for equivalent correlation functions at a range of distances on the network, $\{F_{II}^{(r)}\left(t\right)\}^{r=r_{\text{max}}}_{r=1}$. On the RRG, which lacks a spatial embedding, the notion of Euclidean distance is ill-defined. As such, it is often preferable to consider the shortest path between nodes as a metric of distance. This notion of distance affords the definition of “shells” of nodes surrounding each node on a network, where shells of distance one contain all nodes directly connected to a given node, shells of distance two contain all nodes connected to the nodes of distance one that are not in shells of distance one, etc. (see Fig. 4). This, in turn, means that a set of correlation functions $\{F_{II}^{(\ell)}\left(t\right)\}^{\ell=\ell_{\text{max}}}_{\ell=1}$ can be defined in a manner of Eq. (4), where $\ell$ denotes the shell. In Eqs. (4)-(5), the Euclidean distance $r$ appears only in the enumeration of the correlation function and is not explicitly involved in the system dynamics. As such, Euclidean distance can be exchanged with the shell number $\ell$, where every possible distance from the origin in the lattice is exchanged by the corresponding shell number $\ell$. For example, for the two-dimensional lattice, the distances $r=1,\sqrt{2}/2,\sqrt{3}/2,\dots$, may be exchanged with $\ell=1,2,3,\dots$. In the following sections, we proceed in terms of this notion of shell distance, rather than Euclidean distance.
@@
The dynamics of the pairwise correlation function $F_{II}^{(\ell)}(t)=x_{II}^{(\ell)}/x_{I}^{2}$ can be derived by analyzing the change in both the two-point probability $x_{II}^{(\ell)}$ and the global infected density $x_{I}$. The dynamics of $x_{II}^{(\ell)}$ includes contributions from infection events in which a susceptible node in an IISS pair becomes infected through a third, infected neighbor distinct from the original $I$ node. This process is described by the three–point function $x_{ISI}^{(\ell\ell^{\prime}m)}$, which we approximate using the Kirkwood superposition approximation (KSA). The KSA expresses $x_{ISI}^{(\ell\ell^{\prime}m)}$ as a product of pairwise distributions,
$x_{ISI}^{(\ell\ell^{\prime}m)}\!=\!\frac{x_{SI}^{(\ell)}x_{SI}^{(m)}x_{II}^{(\ell^{\prime})}}{(1\!-\!x_{I})x_{I}^{2}}\!=\!(1-x_{I})x_{I}^{2}F_{SI}^{(\ell)}(t)F_{SI}^{(m)}(t)F_{II}^{(\ell^{\prime})}(t),$ (7)
@@
where $\ell$, $\ell^{\prime}$, and $m$ are the shell distances between nodes 1-2, 1-3, and 2-3 of the triplet, respectively. Since infection occurs by direct contact between a susceptible node and a third infected neighbor, we set $m=1$ and sum over all possible third-party shells $\ell^{\prime}$, weighted by $p(\ell^{\prime}|\ell)$—the conditional probability that a neighbor of a susceptible node in shell $\ell$ lies in shell $\ell^{\prime}$. The total contribution is then equal to $2\beta(1-x_{I})F_{SI}^{(\ell)}(t)\sum_{\ell^{\prime}>0}p(\ell^{\prime}|\ell)\,F_{II}^{(\ell^{\prime})}(t)$, where the factor of two accounts for symmetric SSII and IISS configurations.
@@
Another contribution to the dynamics of $x_{II}^{(\ell)}$ arises from the direct infection of adjacent IISS pairs, which converts them into IIII pairs. This occurs only when $\ell=1$, and can therefore be written as $2\beta\delta_{\ell,1}x_{IS}/x_{I}^{2}=2\beta\delta_{\ell,1}\left(1-x_{I}\right)F_{SI}^{(1)}\left(t\right)/x_{I}$, where again the factor of two follows similarly as before, and $\delta_{n,m}$ is the Kronecker delta. Furthermore, $F_{II}^{(\ell)}(t)$ is also affected by variations in the overall infection density $x_{I}$ that appears in the denominator of $F_{II}^{(\ell)}(t)$. Differentiating $F_{II}^{(\ell)}(t)=x_{II}^{(\ell)}/x_{I}^{2}$ yields a term $2x_{II}^{(\ell)}\dot{x}_{I}/x_{I}^{3}$, which equals $2F_{II}^{(\ell)}(t)\dot{x}_{I}/x_{I}$. Substituting $\dot{x}_{I}$ from Eq. (6) and using the identity $(1-x_{I})F_{SI}^{(\ell)}(t)=1-x_{I}F_{II}^{(\ell)}(t)$ as explained in Eq. (5) recasts this contribution as $2F_{II}^{(\ell)}(t)\,\beta k_{0}(1-x_{I})F_{SI}^{(1)}(t)$. Here, the factor of two reflects the squared normalization in $x_{I}$. Finally, although recovery events reduce both $x_{II}^{(\ell)}$ and $x_{I}$, their effects cancel out in the evolution of $F_{II}^{(\ell)}\left(t\right)$. Recovery of one node in an IIII pair contributes $2F_{II}^{(\ell)}\left(t\right)$ due to pair symmetry, while the change in $x_{I}$ from the same recovery contributes $x_{I}$ to $\dot{x}_{I}$, such that it modifies the denominator in a way that precisely offsets that reduction, resulting in a net zero contribution.
Putting it all together, the dynamics of $F_{II}^{(\ell)}$ are described by the coupled system of ODEs
$\frac{\mathrm{d}F_{II}^{(\ell)}(t)}{\mathrm{d}t}$ $=\left[F_{SI}^{(\ell)}(t)\sum_{\ell^{\prime}>0}p(\ell^{\prime}|\ell)F_{II}^{(\ell^{\prime})}(t)\!+\!\frac{\delta_{\ell,1}}{k_{0}x_{I}}\!-\!F_{II}^{(\ell)}(t)\right]$
@@
where $\ell\in\{1,\dots,\ell_{\max}\}$, and $p(m|n)$ is the conditional probability that a given neighbor of a node in shell $n$ lies in shell $m$. The dynamics can also be written entirely in terms of $F_{II}^{(\ell)}$ by using Eq. (5), which is convenient for numerical solutions.
Notably, Eq. (8) applies both to the regular lattice and to the RRG. For example, in a two‑dimensional square lattice (Fig. 4(a)), a node in $(1,0)$ (first shell) has one neighbor in the zeroth shell, two in the second shell, and one in the third shell, such that $p(k|1)=\delta_{k,0}/4+\delta_{k,2}/2+\delta_{k,3}/4$. Similarly, a node in (1,1)(-1,1) (second shell) has two neighbors in the first shell and two in the fourth shell, giving $p(k|2)=\delta_{k,1}/2+\delta_{k,4}/2$. By translational invariance, the probabilities $p(\ell^{\prime}|\ell)$ are uniform across nodes, and hence are global constants that vary with the lattice dimension.
Figure 4: (a) A two-dimensional regular lattice with its origin at $(0,0)$. The purple, green, blue, and red circles represent the first, second, third, and fourth shells, respectively. (b) RRG shells with node degree $k_{0}=4$. The central node is black, the first shell nodes are colored purple, while the second shell nodes are colored green.
V Computing Shell Probabilities on Regular Random Graphs
To complete the formulation of the dynamics of the pairwise correlation function, Eq. (8), it is left for us to find an explicit expression for $p(\ell^{\prime}|\ell)$ in the case of the RRG. For the latter, the fraction of individuals located in shells beyond distance $\ell$ follows the Gompertz distribution [18]
@@ V Computing Shell Probabilities on Regular Random Graphs
The calculation of the conditional probabilities $p(\ell-1|\ell)$, $p(\ell|\ell)$, and $p(\ell+1|\ell)$ for a target node located in the $\ell$-th shell, is performed iteratively. At each step, these probabilities are determined using the previously computed values for the $(\ell-1)$-th shell and the shell populations, as given by Eq. (10). Specifically, the recursion relation exploits the previously obtained $p(\ell|\ell-1)$, and all the shell populations up to the current $\ell$-th shell, $N_{0},\dots,N_{\ell}$ to advance to the next shell, ensuring that all dependencies are consistently propagated throughout the shell hierarchy. From Eq. (12), it follows that $p\left(\ell-1|\ell\right)=p\left(\ell|\ell-1\right)N_{\ell-1}/N_{\ell}$. Now, the probability $p(\ell|\ell)$ describes the likelihood that a neighbor of a node in shell $\ell$ also lies in the same shell. To compute this, we first condition on the alternative—that the edge does not point back to shell $\ell-1$—which provides the prefactor $\left(1-p(\ell-1|\ell)\right)$. The remaining neighbors of the node can then be distributed either within the same shell $\ell$ or into the next shell $\ell+1$. To evaluate this balance, we count the total number of available “stubs” in shell $\ell$ that can remain internal $k_{0}N_{\ell}$, subtracting those already used to connect back into previous shell $\ell-1$, $k_{0}p\left(\ell-1|\ell\right)N_{\ell}$, and further excluding the trivial self‑connection of the test node $k_{0}\left(1-p\left(\ell-1|\ell\right)\right)$. This yields the numerator of $p(\ell|\ell)$ to be $k_{0}N_{\ell}-k_{0}p\left(\ell-1|\ell\right)N_{\ell}-k_{0}\left(1-p\left(\ell-1|\ell\right)\right)$, reflecting the number of ways that edges can close within shell $\ell$ once the backward connections and self-loops have been removed. The denominator counts the total remaining available “stubs” in the network after all edges to shells $0,\dots,\ell\,$ have been allocated $k_{0}N-\sum_{k=0}^{\ell}k_{0}N_{k}$, plus the corrected contribution from current $\ell$-th shell. After some simplifications, the resulting expression reads
@@
This expression captures the self‑closing probability within the shell as the ratio of “internal same‑shell connections” to “all possible remaining connections”.
@@
The probability $p(\ell+1|\ell)$, representing the likelihood that a neighbor of a node in shell $\ell$ resides in the forward shell $\ell+1$, can be formulated either by enforcing the complementary probability, $p(\ell+1|\ell)=1-p(\ell|\ell)-p(\ell-1|\ell)$, or by employing the same combinatorial logic where the numerator corresponds to available endpoints in the forward shells and the denominator accounts by all remaining stubs. The system is rendered self-consistent by stipulating the boundary condition $p(1|0)=1$, ensuring that the initial shell is reached deterministically from the origin. With this assignment, the recursive relations for the shell populations $N_{k}$ and the transition probabilities $p(k|\ell)$ are fully formulated, thereby defining the complete hierarchy of shell-wise network topology.
To assess the accuracy of the pairwise correlation correction to mean-field theory on the RRG, we directly compare solutions using our analytical framework to numerical simulations of the SIS model. Specifically, we solve a system of $n$ coupled ODEs for the pairwise correlation functions in each shell $\ell$ as given by Eq. (8), with the number of shells $n$ naturally limited so that $\sum_{i=1}^{n}N_{i}\approx N$, and truncated at the shell where $p(n+1|n)<1/N$, indicating that there is fewer than one node in the subsequent shell. The conditional probabilities determine how correlations propagate through the network topology, ensuring consistency in the shell structure. Once solved, the hierarchy of correlation functions, particularly the distance-one correlation function, is then used in Eq. (6) to predict the global infection density. This unified approach enables direct comparison between theory and simulation. We illustrate this on a network of $10^{4}$ nodes, with an initial fraction $10^{-2}$ of infected individuals. The results shown in Fig. 2 show good agreement between the numerical solution of Eq. (8) (blue) and corresponding discrete simulations (red). We note a slight discrepancy between our corrections to mean-field theory and corresponding simulations for infection rates $\beta$ close to the critical values $\beta_{c}k_{0}/\gamma=1.33$ and $1.20$ for $k_{0}=4$ and $k_{0}=6$, respectively, according to the formula $\beta_{c}k_{0}/\gamma=k_{0}/(k_{0}-1)$ for homogeneous networks [19]. These critical values separate the regime where the infection persists in the population from the regime in which the epidemic dies out deterministically. The discrepancies near these critical points, where epidemic dynamics are dominated by strong fluctuations and long-range correlations, are inherently difficult to attenuate in low-order moment closure, such as the KSA, and have been discussed in previous works [11, 12]. In general, the agreement between model simulations and numerical solutions using our modified theory improves with the increase of the node degree $k_{0}$ or as the infection rate $\beta$ increases.
VI A modified steady state analysis on regular random networks.

### v2

@@ III Dependence of nearest-neighbor correlation on network topology
We now further motivate the development of bespoke corrections to mean-field theory for the RRG by using numerical simulations to highlight significant variations in $F_{II}^{(1)}$, and hence global infection dynamics, with the underlying topology of regular networks (i.e., networks with a fixed node degree $k_{0}$). To do this, we investigate the effects of rewiring by taking initially two- and three-dimensional regular lattices, which possess a clear spatial embedding and strict local neighborhoods, and randomly exchange their edges with probability $p$. This approach, inspired by the methodology of Watts and Strogatz [21], constructs networks that preserve the fixed node degree but become increasingly random, progressively destroying the original spatial structure as the probability of two spatially adjacent nodes remaining connected decreases with $p$. The edge swapping algorithm we use is a modification of an algorithm from Wang et al. [20] and is presented in Appendix A. It yields randomized network structures that converge to the RRG as $p$ becomes ${\cal O}(1)$.
@@
To study how increasing topological randomness affects spatial correlations between adjacent nodes and infection trajectories, we perform a parameter sweep over the per-node infection rate, $k_{0}\beta$, and the edge-swapping probability, $p$, on two- and three-dimensional square lattices, with $k_{0}=4$ and $k_{0}=6$, respectively. We find that increasing either $p$ (thereby reducing the mean shortest path length) or $k_{0}\beta$ substantially suppresses spatial correlations. This is reflected in shorter relaxation times (Figs. 3(a) and 3(c)), defined as $\tau_{\mathrm{relax}}=t_{\mathrm{th}}-t_{\mathrm{peak}}$. Here, $t_{\mathrm{peak}}$ is the time at which $F^{(1)}_{II}$ reaches its maximum, $F^{(1)}_{II,\mathrm{peak}}$, $t_{\mathrm{th}}$ is the time at which it decays to within 10% of its steady-state value, i.e., $F^{(1)}_{II}(t_{\mathrm{th}})=\bar{F}^{(1)}_{II}+0.1(F^{(1)}_{II,\mathrm{peak}}-\bar{F}^{(1)}_{II})$, while $\bar{F}^{(1)}_{II}$ denotes the steady-state value estimated from the long-time tail. The suppression is also evident in the reduced steady-state values of $F_{II}^{(1)}$ (Figs. 3(b) and 3(d)). Overall, adding topological randomness increases the steady-state infection density and accelerates the dynamics, making its effect qualitatively similar to that of increasing the infection rate $\beta$. Similar sensitivities to network topology have been reported in the proportion of recovered individuals in related epidemic models [5, 7].
Notably, the results in Fig. 3 suggest that the corrections to mean-field theory proposed for the regular lattice, where $p=0$ [1, 11], are not appropriate for RRGs, where $p={\cal O}(1)$, because of the strong dependence of $F_{II}^{(1)}$ on $p$. Below, we therefore derive a correction to mean-field theory based on random network theory, which outperforms both the standard mean-field theory and SPM.
@@
Figure 3: Relaxation time and steady-state values for the SIS process versus $k_{0}\beta$ and $p$. Panels (a) and (c) show the natural logarithm of the relaxation times, defined as the time between the peak ($t_{\mathrm{peak}}$) and the first subsequent time $t_{\mathrm{th}}$ at which $F^{(1)}_{II}(t)$ has decayed to $0.1$ above its steady-state value $\bar{F}^{(1)}_{II}$, whilst (b) and (d) show $\ln(\bar{F}^{(1)}_{II}-1)$ for $k_{0}=6$ and $4$, respectively. This parameter sweep shows that increasing $p$ and increasing $k_{0}\beta$ both reduce the steady-state value of $F^{(1)}_{II}$ and the relaxation time of the process. Simulations were executed on networks with $N\sim 10^{4}$ nodes (with the exact value depending on the lattice dimension) and $\gamma=1$. In all simulations, $0.01N$ infectious individuals are randomly placed on the network, with all other nodes occupied by susceptible individuals, and results are averaged over 100 simulations.
@@ IV Dynamics of the Pairwise Correlation Function
We now formulate generalized dynamics for the distance-one pairwise correlation function, $F_{II}^{(1)}\left(t\right)$, for both the lattice and the RRG, using a series of hierarchical ODEs for equivalent correlation functions at a range of distances on the network, $\{F_{II}^{(r)}\left(t\right)\}^{r=r_{\text{max}}}_{r=1}$. On the RRG, which lacks a spatial embedding, the notion of Euclidean distance is ill-defined. As such, it is often preferable to consider the shortest path between nodes as a metric of distance. This notion of distance affords the definition of “shells” of nodes surrounding each node on a network, where shells of distance one contain all nodes directly connected to a given node, shells of distance two contain all nodes connected to the nodes of distance one that are not in shells of distance one, etc. (see Fig. 4). This, in turn, means that a set of correlation functions $\{F_{II}^{(\ell)}\left(t\right)\}^{\ell=\ell_{\text{max}}}_{\ell=1}$ can be defined in a manner of Eq. (4), where $\ell$ denotes the shell, and $\ell_{\max}$ is the maximum shell index, defined explicitly in Sec. V. In Eqs. (4)-(5), the Euclidean distance $r$ appears only in the enumeration of the correlation function and is not explicitly involved in the system dynamics. Thus, $r$ can be seamlessly replaced by the shell index $\ell$; e.g., on the two-dimensional lattice, the distances $r=1,\sqrt{2}/2,\sqrt{3}/2,\dots$, map directly to $\ell=1,2,3,\dots$ (see Fig. 4). Henceforth, we formulate the dynamics entirely in terms of this shell index.
@@
The dynamics of the pairwise correlation function $F_{II}^{(\ell)}=x_{II}^{(\ell)}/x_{I}^{2}$ can be derived by analyzing the change in both the two-point probability $x_{II}^{(\ell)}$ and global infected density $x_{I}$. The former includes contributions from infection events in which a susceptible node in an II-$S$ pair becomes infected through a third, infected neighbor distinct from the original $I$ node. This process is described by the three-point function $x_{ISI}^{(\ell\ell^{\prime}m)}$, evaluated via the Kirkwood superposition approximation (KSA). The KSA expresses $x_{ISI}^{(\ell\ell^{\prime}m)}$ as a product of pairwise distributions,
$x_{ISI}^{(\ell\ell^{\prime}m)}\!=\!\frac{x_{SI}^{(\ell)}x_{SI}^{(m)}x_{II}^{(\ell^{\prime})}}{(1\!-\!x_{I})x_{I}^{2}}\!=\!(1-x_{I})x_{I}^{2}F_{SI}^{(\ell)}F_{SI}^{(m)}F_{II}^{(\ell^{\prime})},$ (7)
where $\ell$, $\ell^{\prime}$, and $m$ are the shell distances between nodes 1-2, 1-3, and 2-3 of the triplet, respectively. This closure differs from that used in the SPM, where the triplet closure is applied only to nearest-neighbor triples (with edges connecting nodes 1-2 and 2-3, yielding $m=\ell=1$), and correlations beyond nearest neighbors are ignored by setting $F_{II}^{(\ell^{\prime})}\!\!=\!x_{II}^{(\ell^{\prime})}/x_{I}^{2}\!=\!1$ for all $\ell^{\prime}>1$. Thus, Eq. (7) is retained only when there is an edge connecting nodes 1-3 ($\ell^{\prime}=1$), forming a triangle 1-2-3. Yet, on both the square lattice and RRG, the probability of such triangles vanishes for network size $N\gg 1$. Hence, the SPM reduces to a closure $x_{ISI}=x_{SI}^{2}/(1-x_{I})$, neglecting longer-range spatial correlations.
@@
We now outline the derivation of the dynamics of $F_{II}^{(\ell)}$. To account for this third-party infection process, we set $m=1$ and sum over all possible third-party shells $\ell^{\prime}$, weighted by $p(\ell^{\prime}|\ell)$, the conditional probability that a neighbor of a susceptible node in shell $\ell$ lies in shell $\ell^{\prime}$. The total contribution to $\mathrm{d}F_{II}^{(\ell)}/\mathrm{d}t$ is then equal to $2\beta(1-x_{I})F_{SI}^{(\ell)}(t)\sum_{\ell^{\prime}>0}k_{0}p(\ell^{\prime}|\ell)\,F_{II}^{(\ell^{\prime})}(t)$, where the factor of two accounts for symmetric $S$-II and II-$S$ pairs.
@@
Another contribution to the dynamics of $x_{II}^{(\ell)}$ arises from the direct infection of adjacent II-$S$ pairs, which converts them into II-II pairs. This occurs only when $\ell=1$, and can thus be written as $2\beta\delta_{\ell,1}x_{SI}^{(1)}/x_{I}^{2}=2\beta\delta_{\ell,1}\left(1-x_{I}\right)F_{SI}^{(1)}/x_{I}$, where the factor of two follows similarly as before, and $\delta_{n,m}$ is the Kronecker delta. Furthermore, $F_{II}^{(\ell)}$ is also affected by variations in the overall infection density $x_{I}$ that appear in the denominator of $F_{II}^{(\ell)}$. Differentiating $F_{II}^{(\ell)}=x_{II}^{(\ell)}/x_{I}^{2}$ yields a term $2x_{II}^{(\ell)}(\mathrm{d}x_{I}/\mathrm{d}t)/x_{I}^{3}$, which equals $2F_{II}^{(\ell)}(\mathrm{d}x_{I}/\mathrm{d}t)/x_{I}$. Substituting the infectious term from Eq. (6) and using the identity $(1-x_{I})F_{SI}^{(\ell)}=1-x_{I}F_{II}^{(\ell)}$ from Eq. (5) recasts this contribution as $2F_{II}^{(\ell)}\,\beta k_{0}(1-x_{I})F_{SI}^{(1)}$. Here, the factor of two reflects the squared normalization in $x_{I}$.
Finally, while recovery events reduce both $x_{II}^{(\ell)}$ and $x_{I}$, their effects cancel out in the evolution of $F_{II}^{(\ell)}\left(t\right)$. Specifically, recovery of either node in an II-II pair decreases the numerator $x_{II}^{(\ell)}$ at a rate $2x_{II}^{(\ell)}$, contributing $2F_{II}^{(\ell)}$ to the time derivative. Simultaneously, the recovery term in Eq. (6) acts through the derivative of the $x_{I}^{2}$ denominator, $2F_{II}^{(\ell)}(\mathrm{d}x_{I}/\mathrm{d}t)/x_{I}$, yielding an exactly compensating $2F_{II}^{(\ell)}$ term, resulting in a net zero effect.
In total, the dynamics of $F_{II}^{(\ell)}$ are described by the coupled system of ODEs
$\frac{\mathrm{d}F_{II}^{(\ell)}}{\mathrm{d}t}$ $=\left[F_{SI}^{(\ell)}\sum_{\ell^{\prime}>0}p(\ell^{\prime}|\ell)F_{II}^{(\ell^{\prime})}\!+\!\frac{\delta_{\ell,1}}{k_{0}x_{I}}\!-\!F_{II}^{(\ell)}\right]$
@@
where $\ell\in\{1,\dots,\ell_{\max}\}$, and $p(m|n)$ is the conditional probability that a given neighbor of a node in shell $n$ lies in shell $m$. Note that, the dynamics can also be written entirely in terms of $F_{II}^{(\ell)}$ by using Eq. (5), which is convenient for numerical solutions.
Importantly, Eq. (8) applies both to the regular lattice and to the RRG. For example, in a two‑dimensional square lattice (Fig. 4(a)), a node in $(1,0)$ (the first shell) has one neighbor in the zeroth shell, two in the second shell, and one in the third shell, such that $p(k|1)=\delta_{k,0}/4+\delta_{k,2}/2+\delta_{k,3}/4$. Similarly, a node in (1,1)(-1,1) (the second shell) has two neighbors in the first shell and two in the fourth shell, giving $p(k|2)=\delta_{k,1}/2+\delta_{k,4}/2$. By translational invariance, the probabilities $p(\ell^{\prime}|\ell)$ are uniform across nodes, and hence are global constants that vary with the lattice dimension.
Figure 4: (a) A two-dimensional regular lattice with its origin at $(0,0)$. The purple, green, blue, and red circles represent the first, second, third, and fourth shells, respectively. (b) RRG shells with node degree $k_{0}=4$. The central node is black, and the shells follow the same color scheme as in (a).
V Computing Shell Probabilities on Random Regular Graphs
To complete the formulation of the dynamics of the pairwise correlation function, Eq. (8), it remains to find an explicit expression for $p(\ell^{\prime}|\ell)$ in the case of the RRG. For the latter, the fraction of individuals located in shells beyond distance $\ell$ follows the Gompertz distribution [13]
@@ V Computing Shell Probabilities on Random Regular Graphs
The calculation of the conditional probabilities $p(\ell-1|\ell)$, $p(\ell|\ell)$, and $p(\ell+1|\ell)$ for a target node in the $\ell$-th shell is performed iteratively. At each step, the recursion exploits the previously obtained $p(\ell|\ell-1)$ and the shell populations $N_{0},\dots,N_{\ell}$ to advance to the next shell. From Bayes’ theorem (Eq. (12)), the backward probability is simply $p\left(\ell-1|\ell\right)=p\left(\ell|\ell-1\right)N_{\ell-1}/N_{\ell}$. Next, the probability $p(\ell|\ell)$ describes the likelihood that a neighbor of a node in shell $\ell$ also lies in the same shell. We compute this using a combinatorial stub-matching argument. The number of available “stubs” (half-edges yet to be paired to form edges) in shell $\ell$ that can connect internally is the total capacity $k_{0}N_{\ell}$, minus the stubs already paired to form edges to the previous shell $\ell-1$, and excluding trivial self-connections. Dividing this by the total remaining available stubs in the network (after accounting for all connections in shells $0,\dots,\ell$) and simplifying the resulting algebraic expression, yields:
@@
This expression captures the self-closing probability within the shell as the ratio of available internal same-shell connections to all possible remaining connections.
@@
The quantity $p(\ell+1|\ell)$, representing the probability that a neighbor of a node in shell $\ell$ resides in the forward shell $\ell+1$, can be formulated either by enforcing the complementary probability, $p(\ell+1|\ell)=1-p(\ell|\ell)-p(\ell-1|\ell)$, or by employing the same combinatorial logic where the numerator corresponds to available endpoints in the forward shells and the denominator accounts for all remaining stubs. The system is rendered self-consistent by stipulating the boundary condition $p(1|0)=1$, ensuring that the initial shell is reached deterministically from the origin. With this assignment, the recursive relations for the shell populations $N_{k}$ and the transition probabilities $p(k|\ell)$ are fully formulated, thereby defining the complete hierarchy of shell-wise network topology.
To assess the accuracy of our MPM against the SPM and mean-field theories on the RRG, we directly compare analytical solutions to numerical simulations of the SIS model. Specifically, we solve a system of coupled ODEs for the pairwise correlation functions in each shell $\ell$ (Eq. (8)), truncated at a maximum shell index $\ell_{\max}$. We dynamically define $\ell_{\max}$ as the smallest shell index for which the expected number of nodes in the next shell falls below unity ($N_{\ell_{\max}+1}<1$), determined via Eq. (10). This ensures that shells beyond $\ell_{\max}$ contain a negligible fraction of the network and thus, contribute minimally to the correlation dynamics. For the network sizes considered ($N=10^{4}$), this yields $\ell_{\max}\approx 5$99 for $k_{0}=3$1010 (where $\ell_{\max}$ decreases with increasing $k_{0}$), capturing the essential spatial structure while remaining computationally tractable. The conditional probabilities determine how correlations propagate through the network topology, ensuring consistency in the shell structure. Once solved, the hierarchy of correlation functions, particularly the distance-one correlation function, is then used in Eq. (6) to predict the global infection density.
This unified approach enables direct comparison between theory and simulation. We illustrate this on a network of $N=10^{4}$ nodes, with an initial fraction $x_{I}(0)=10^{-2}$ of infected individuals. The results shown in Fig. 2 display good agreement between the numerical solution of Eq. (8) (blue) and discrete simulations (red). Notably, the MPM systematically outperforms the SPM (dark green), which overestimates the steady-state infection density and underestimates the correlation function $F_{II}^{(1)}$, especially at lower node degrees. Note that, the slight discrepancy between the MPM and simulations is due to the proximity of $\beta$ to the critical infection rate: $\beta_{c}k_{0}/\gamma=k_{0}/(k_{0}-1)$ for the RRG [12]; the latter separates the endemic regime, $\beta>\beta_{c}$ (persistence of infection), from the absorbing regime, $\beta<\beta_{c}$ (epidemic dies out deterministically). Close to the critical points ($\beta_{c}k_{0}/\gamma=1.5$ and $1.33$ for $k_{0}=3$ and $4$, respectively) the epidemic dynamics are dominated by strong fluctuations and long-range correlations, and are inherently more difficult to capture in low-order moment closure, such as the KSA [1, 11]. As was shown (see also below), the agreement between model simulations and our MPM improves with increasing node degree $k_{0}$ or infection rate $\beta$.
VI A modified steady-state analysis on random regular networks